2 5/6= t/3
solve for t
need to solve using reciprocal

Answer by checkley77(12844)

2 5/6= t/3 solve for t
(2*6+5)/6=t/3
(12+5)/6=t/3
17/6=t/3 multiply by 6/17
1=t/3*6/17
1=6t/51 multiply by 51/6
t=51/6
t=51/6
t=8.5 ans.